# Unexpected but informative: What fixation-related potentials tell us about the processing of confusing program code

### v2

@@ 4.5.5 EEG preprocessing
After this preparation, we synchronized the event triggers in the EEG files with the behavioral data. For two trials in the behavioral data, we were not able to assign the corresponding EEG annotations due to hardware malfunctions. Then, we cut each participant’s continuous EEG file into individual trials (from start of the fixation cross to end of the snippet presentation), resulting in trials usable for ERP analysis. For the FRP analysis, this leaves us with () trials due to an overlap of one trial in the excluded trials within the EEG and the eye-tracking data preparation.
@@ 4.6.1 Behavioral analysis
We analyzed the behavioral data using linear mixed effects models. As a model family, we chose a Gaussian error model with logarithmic values for fixation duration. Answer correctness and subjective difficulty rating were coded as binary variables (, and , , respectively) and analyzed using binomial mixed effects models. For each behavioral measure, the regression models included the fixed parameter condition to identify whether a trial was confusing or clean (Condition: , ), as well as the block number (BlockNo: –) and the trial number within a block (ItemOrder: –) as well as their interactions to account for potential learning and habituation effects. Linear mixed effects models included random intercepts and random slopes for condition under participants and items; the fixed effects structure was reduced using backward model selection based on the Akaike information criterion and a significance threshold of .
@@ 4.6.2 FRP analysis
To calculate the FRPs, we truncated the EEG segment of each trial relative to the start of the fixation into the epochs from to including a baseline from to . The offset to stimulus onset leads to a further reduction of trials, since fixations at the end of the trial might not have the entire epoch length available within the snippet view and thus cannot be included, leaving us with () trials. Next, the epochs were baseline-corrected to reduce the influence of pre-stimulus activity. That is, for each electrode and each sample point within the epoch ( to ), the average pre-stimulus voltage was subtracted from the EEG signal. For determining trials containing artifacts, we ignored all EOG electrodes as well as the reference electrodes. We excluded all epochs that violate one of the following voltage constraints on any of the electrodes, which indicates the presence of an artifact in the epoch: First, a voltage step between two consecutive samples must be lower than µV. Second, the maximal voltage difference within should not exceed µV. Last, the absolute amplitude after baseline correction must remain belowµV at any point in time within the epoch. We excluded trials ( remaining). As a result, on average, (; range: –) trials per condition and participant are suitable for the FRP analysis. Finally, we averaged the epochs per participant, condition, and electrode.
For statistical analysis, we calculated the FRP () difference wave per participant and electrode to use it as a dependent variable in a pairwise cluster permutation test [65]. This test identifies possibly significant data points consisting of time points and electrodes, in which the test statistics exceed the significance threshold () of a single paired t-test, and clusters these if, at least, two spatially or temporally adjacent points reached the significance threshold. Spatial adjacency was defined by adjacent electrodes (e.g., Fp1 and Fp2, see Figure 3a) and temporal adjacency was defined by two consecutive time points. Then, the test evaluates the significance of each detected cluster by a permutation test, which has the advantage of inherently controlling for multiple testing [66]. To maximize the power in the permutation test, we followed the suggestions of Luck et al. [24]: The epochs were down-sampled to 100 Hz, and electrodes T7, T8, O1, and O2 were not included in the analysis, since we did not expect any significant effects at these electrode sites.
Additionally, we reduced the analysis interval to to . We performed two pairwise cluster permutation tests with this input for the difference being both a significantly smaller () and a significantly higher () amplitude than µV and consider all resulting clusters with an alpha level of as significant. For each of the permutation tests, we performed the maximum number of permutations () [66].
@@ Acknowledgements
We would like to thank all the participants of the study, who made this research possible, as well as our research assistants Youssef Abdelsalam, Yannick Lehmen, and Marcellius William Suntoro, who helped with the technical setup.
@@
This work has been supported by the European Union as part of the ERC Advanced Grant Brains On Code (101052182) and by the German Research Foundation as part of Transregional Collaborative Research Centre 248 Foundations of Perspicuous Software Systems (389792660). Axel Mecklinger and Regine Bader were funded by the Deutsche Forschungsgemeinschaft (DFG) – Project ID 232722074 CRC 1102, project A6.
@@
This experiment used code available in Python (v. 3.11.5), as well as R (v. 4.3.2) for linear mixed effect models. Additionally, we used other open-source (PsychoPy, v. 2021.2.3) and commercial applications (Tobii Eye-Tracker Manager (v. 2.6.0), BrainVision Recorder (v. 1.20.0801), and BrainVision Analyzer (v. 2.3.0.8300)) to perform the experiment. We deposited the scripts and content in a GitHub Repository https://github.com/brains-on-code/AoC-FRP-Code and they are available for unrestricted open access under a CC-BY license.

### v3

@@ 4.5.5 EEG preprocessing
After this preparation, we synchronized the event triggers in the EEG files with the behavioral data. For two trials in the behavioral data, we were unable to assign the corresponding EEG annotations due to hardware malfunctions. Then, we cut each participant’s continuous EEG file into individual trials (from the start of the fixation cross to the end of the snippet presentation), resulting in trials usable for ERP analysis. For the FRP analysis, this leaves us with () trials due to an overlap of one trial in the excluded trials within the EEG and the eye-tracking data preparation.
@@ 4.6.1 Behavioral analysis
We analyzed the behavioral data using linear mixed effects models. As a model family, we chose a Gaussian error model with logarithmic values for fixation duration. Answer correctness and subjective difficulty ratings were coded as binary variables (, and , , respectively) and analyzed using binomial mixed effects models. For each behavioral measure, the regression models included the fixed parameter condition to identify whether a trial was confusing or clean (Condition: , ), as well as the block number (BlockNo: –) and the trial number within a block (ItemOrder: –) along with their interactions to account for potential learning and habituation effects. Linear mixed effects models included random intercepts and random slopes for condition under participants and items; the fixed effects structure was reduced using backward model selection based on the Akaike information criterion and a significance threshold of .
@@ 4.6.2 FRP analysis
To calculate the FRPs, we truncated the EEG segment of each trial relative to the start of the fixation into the epochs from to including a baseline from to . The offset of the fixation compared to stimulus onset leads to a further reduction of trials since fixations at the end of the trial might not have the entire epoch length available within the snippet view and thus cannot be included, leaving us with () trials. Next, the epochs were baseline-corrected to reduce the influence of pre-stimulus activity. That is, for each electrode and each sample point within the epoch ( to ), the average pre-stimulus voltage was subtracted from the EEG signal. For determining trials containing artifacts, we ignored all EOG electrodes as well as the reference electrodes. We excluded all epochs that violated one of the following voltage constraints on any of the electrodes, which indicates the presence of an artifact in the epoch: First, a voltage step between two consecutive samples must be lower than µV. Second, the maximal voltage difference within should not exceed µV. Last, the absolute amplitude after baseline correction must remain belowµV at any point in time within the epoch. We excluded trials ( remaining). As a result, on average, (; range: –) trials per condition and participant are suitable for the FRP analysis. Finally, we averaged the epochs per participant, condition, and electrode.
For statistical analysis, we calculated the FRP () difference wave per participant and electrode to use it as a dependent variable in a pairwise cluster permutation test [69]. This test identifies possibly significant data points consisting of time points and electrodes in which the test statistics exceed the significance threshold () of a single paired t-test, and clusters these if at least two spatially or temporally adjacent points reach the significance threshold. Spatial adjacency was defined by adjacent electrodes (e.g., Fp1 and Fp2, see Figure 3a) and temporal adjacency was defined by two consecutive time points. Then, the test evaluates the significance of each detected cluster by a permutation test, which has the advantage of inherently controlling for multiple testing [70]. To maximize the power in the permutation test, we followed the suggestions of Luck et al. [24]: The epochs were down-sampled to 100 Hz, and electrodes T7, T8, O1, and O2 were not included in the analysis since we did not expect any significant effects at these electrode sites.
Additionally, we reduced the analysis interval to to . We performed two pairwise cluster permutation tests with this input for the difference being both a significantly smaller () and a significantly higher () amplitude than µV and considered all resulting clusters with an alpha level of as significant. For each of the permutation tests, we performed the maximum number of permutations () [70].
@@ Acknowledgements
We would like to thank all the participants of the study who made this research possible, as well as our research assistants, Youssef Abdelsalam, Yannick Lehmen, and Marcellius William Suntoro, who helped with the technical setup.
@@
This work has been supported by the European Union as part of the ERC Advanced Grant Brains On Code (101052182) and by the German Research Foundation as part of the Transregional Collaborative Research Center 248 Foundations of Perspicuous Software Systems (389792660). Axel Mecklinger and Regine Bader were funded by the Deutsche Forschungsgemeinschaft (DFG) – Project ID 232722074 CRC 1102, project A6.
@@
This experiment used code available in Python (v3.11.5) as well as R (v4.3.2) for linear mixed effects models. Additionally, we used other open-source (PsychoPy, v2021.2.3) and commercial applications (Tobii Eye-Tracker Manager (v2.6.0), BrainVision Recorder (v1.20.0801), and BrainVision Analyzer (v2.3.0.8300)) to perform the experiment. We deposited the scripts and content in a GitHub Repository https://github.com/brains-on-code/AoC-FRP-Code and they are available for unrestricted open access under a CC-BY license.